### 16 Mar 2026 v1

@@ Appendix A ALMA archival data reduction
The 1.3 mm data from project 2016.1.00039.S (P.I. Dunham) were obtained on November 1st 2017, only 3 weeks before our 0.89 mm observations. The data were obtained with 49 antennas of the 12-array in single field dual-polarization mode, dedicating one spectral window of 1.875 GHz bandwidth to observe continuum and five more spectral windows to observe spectra line emission, comprising an aggregate bandwidth of 2.262 GHz. The average precipitable water-vapor column of the observations was 0.5 mm. The baselines ranged from 113 m to 13.9 km. The maximum recoverable scale was 05. Bandpass and flux calibrations were made using QSO J0510+1800, while QSO J0431+1731 was used as a phase calibrator. The science time on-source was 31 minutes. Data were calibrated using the CASA pipeline, version 5.1.1. Continuum imaging was performed using the task tclean in multi-frequency synthesis mode and a multi-scale clean deconvolution algorithm with scales of 0, 004, and 01. A cellsize of 0004 was used, with Briggs weighting and a robust parameter = 0.5, providing a synthesized beam of 004003 and PA of -36°. We tried the self-calibration of the data, but it did not work due to the faintness of the source. The uvrange of the data spans from 53 to 9200.
@@ Appendix B ALMA Band 7 data
To further investigate the presence of two compact sources in the ALMA data, we have compared different data reductions with different robust parameters and a cut in uvrange 200 . We have explored robust of 0.5, 0.3, -0.25, -0.5 and -1, and we display the images with -0.25 and 0.5 in Figs. 1 and 5, respectively. The left panel of Fig. 5 shows the 0.89 mm image with robust 0.5, in which we detect a central source resolved into two bright peaks separated by 004, and an additional extended emission towards the north-west, in the direction where we detect the secondary component in the 1.3 mm data. As the robust is decreased, giving more weight to the longest baselines, the two central peaks set apart, and show a separation of 0054 for a robust of -0.25, as shown in Section 2.1). Although this particular robust allows us to resolve the binary, it has the disadvantage of increasing the image noise ( 6.6e-5 Jy/beam) in comparison with the robust of 0.5 (4.9e-5 Jy/beam).
@@
Due to the fact that a negative robust enhances artificial clumpy emission, we have also generated a 0.89 mm self-calibrated image with robust of 0.5 and a cut in uvrange of uvrange1100 , giving more weight to the most extended baselines. The result is displayed in Fig. 6 (left panel). The image with the uv cut shows a strong source to the east plus an additional well-separated source to the west at a separation of 80 mas. The two bright emissions are connected by a stream of material. For comparison, we have also applied a cut in uvrange of 1100 to the 1.3 mm image (6, right panel), recovering the reported binary with the detection of two compact sources at a separation of 80 mas.
Appendix C ALMA parameters of the sources detected at 0.89 mm and 1.3 mm continuum images.
Table 1 includes the coordinates, peak intensities, fluxes, and deconvolved sizes of the two sources detected in the 0.89 mm and 1.3 mm ALMA images. Note that, for the reasons explained above, the 0.89 mm measurements have been performed both in the self-calibrated image shown in Figure 1, and in the non self-calibrated image displayed in the right panel of Figure 5. The ALMA flux calibration accuracy is 10 in the two bands (see e.g. 10.5281/zenodo.4511522). Positions and peak intensities of the two sources have been derived using imstat. A double Gaussian has been fitted to the most compact emission of the data to estimate the flux density and the deconvolved size of the two binary components in the two bands.
@@ Appendix E VLA observations of IRAM04191
Notes: for K and Ka, the measurements are obtained on the images with natural weighting, while for Q they are measured on the images with natural weighting and a taper of 2000 ; 1 We provide the size in and the position angle (PA) in degrees, with their associated uncertainties; 2 We provide the beam size and its position angle (PA).

### 19 Mar 2026 v2

@@ Acknowledgements.
- Substellar candidates at the earliest stages: The SUCANES database. A&A 695, pp. A13. External Links: Document, 2412.09091, ADS entry Cited by: Figure 4.
@@ Appendix A ALMA archival data reduction
The 1.3 mm data from project 2016.1.00039.S (P.I. Dunham) were obtained on November 1, 2017, only about three weeks before our 0.89 mm observations. The data were obtained with 49 antennas of the 12-array in single field dual-polarization mode, dedicating one spectral window of 1.875 GHz bandwidth to observe continuum and five more spectral windows to observe spectra line emission, comprising an aggregate bandwidth of 2.262 GHz. The average precipitable water-vapor column of the observations was 0.5 mm. The baselines ranged from 113 m to 13.9 km. The maximum recoverable scale was 05. Bandpass and flux calibrations were made using QSO J0510+1800, while QSO J0431+1731 was used as a phase calibrator. The science time on-source was 31 minutes. Data were calibrated using the CASA pipeline, version 5.1.1. Continuum imaging was performed using the task tclean in multi-frequency synthesis mode and a multi-scale clean deconvolution algorithm with scales of 0, 004, and 01. A cellsize of 0004 was used, with Briggs weighting and a robust parameter = 0.5, providing a synthesized beam of 004003 and PA of -36°. We tried the self-calibration of the data, but it did not work due to the faintness of the source. The uvrange of the data spans from 53 to 9200.
@@ Appendix B ALMA Band 7 data
To further investigate the presence of two compact sources in the ALMA data, we compared images generated using different robust and a cut in uvrange 200 . We have explored robust of 0.5, 0.3, -0.25, -0.5 and -1, and we display the images with -0.25 and 0.5 in Figs. 1 and 5, respectively. The left panel of Fig. 5 shows the 0.89 mm image with robust 0.5, in which we detect a central source resolved into two bright peaks separated by 004, and an additional extended emission toward the northwest, in the direction where we detect the secondary component in the 1.3 mm data. As the robust is decreased, giving more weight to the longest baselines, the two central peaks set apart, and show a separation of 0054 for a robust of -0.25, as shown in Section 2.1. Although this particular robust allows us to resolve the binary, it has the disadvantage of increasing the image noise ( 6.6e-5 Jy/beam) in comparison with the robust of 0.5 (4.9e-5 Jy/beam).
@@
Due to the fact that a negative robust enhances artificial clumpy emission, we have also generated a 0.89 mm self-calibrated image with robust of 0.5 and a cut in range of 1100 , giving more weight to the most extended baselines. The result is displayed in Fig. 6 (left panel). The image with the cut shows a strong source to the east plus an additional well-separated source to the west at a separation of 80 mas. The two bright emissions are connected by a stream of material. For comparison, we have also applied a cut in range of 1100 to the 1.3 mm image (6, right panel), recovering the reported binary with the detection of two compact sources at a separation of 80 mas.
Appendix C ALMA parameters of the sources detected at 0.89 mm and 1.3 mm continuum images
Table 1 includes the coordinates, peak intensities, fluxes, and deconvolved sizes of the two sources detected in the 0.89 mm and 1.3 mm ALMA images. Note that, for the reasons explained above, the 0.89 mm measurements have been performed both in the self-calibrated image shown in Figure 1, and in the non self-calibrated image displayed in the right panel of Figure 5. The ALMA flux calibration accuracy is 10 in the two bands (see, e.g., 10.5281/zenodo.4511522). Positions and peak intensities of the two sources have been derived using imstat. A double Gaussian has been fit to the most compact emission of the data to estimate the flux density and the deconvolved size of the two binary components in the two bands.
@@ Appendix E VLA observations of IRAM04191
Notes: For K and Ka, the measurements were obtained on the images with natural weighting, while for Q they were measured on the images with natural weighting and a taper of 2000 . 1 We provide the size in milliarcseconds and the PA in degrees, with their associated uncertainties. 2 We provide the beam size and its PA in each band.